OpenStudy (anonymous):

I'm trying to find the oredered pair for 2x-y=8 and 9x-4=37

hartnn (hartnn):

hello, @x95b20 \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \) can you check the 2nd equation ? is it 9x-4=37 or 9x-4y=37 or something else ?

OpenStudy (anonymous):

it is 9x-4=37

hartnn (hartnn):

if thats the case, then you can directly find 'x' from that equation! first add 4 on both sides, what do u get ?

OpenStudy (anonymous):

9x=41

OpenStudy (anonymous):

so would x=4.5

hartnn (hartnn):

yes, so x =41/9 i still think its 9x-4y=37 maybe there's typing mistake in your question ?

OpenStudy (anonymous):

you're right! my typo...sorry

hartnn (hartnn):

9x=41 x=41/9 = 4.555 is not an integer solution whereas with 9x-4y=37 we get integers ,,,,good! no problem :)

hartnn (hartnn):

so, do u know about elimination method ? or substitution method ??

OpenStudy (anonymous):

am trying to learn the substitution method..

hartnn (hartnn):

good, so can you isolate 'y' from 2x-y =8 ?? y=.... ?

OpenStudy (anonymous):

I think I have to get rid of the 2... So would I subract 2 from each side?

hartnn (hartnn):

no, to isolate 'y' , you need to get rid of entire 2x term with y so, easier thing to do will be add 'y' to both sides first and then subtract 8 from both sides....try ?

OpenStudy (anonymous):

2x-y+y=8 2x=16?

OpenStudy (anonymous):

2x=0 because the subraction got rid of the 8

hartnn (hartnn):

oh, but when you add 'y' to BOTH SIDES, you add it to RIGHT side too! so, 2x-y +y = 8+ y gives 2x = 8+y got this ?? now you can subtract 8 from BOTH sides...

OpenStudy (anonymous):

yes.. now I need to subract 8 from each side...

hartnn (hartnn):

correct, do that and tell me what u get ?

OpenStudy (anonymous):

-6x=0y

hartnn (hartnn):

oh, you CANNOT combine 2x and -8 to get -6x or anything, they are NOT like terms...only like terms can be combined! so, 2x-8 = 8-8+y gives 2x-8 = 0+y or y= 2x-8 clear till here ?

OpenStudy (anonymous):

since 8 doesn't have an x I cannot subract it then

hartnn (hartnn):

correct, 8 doesn't have x but you can subtract 8, just cannot combine it with 2x there, because as you said, 8 doesn't have x

OpenStudy (anonymous):

ok

hartnn (hartnn):

so, now can we go further with the problem?

OpenStudy (anonymous):

we can divide it by 2 so x=4?

hartnn (hartnn):

in y= 2x-8 ? there's still a 'y' there... thats where 2nd equation comes into picture! :) we got y as 2x-8, now we plug this in 2nd equation for 'y' 9x-4y=37 9x-4 (2x-8) =37 got this step ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Im plugging in the first y where this y is

hartnn (hartnn):

now there's only 'x' in the equation 9x-4 (2x-8) =37 can you get the value of 'x' ?

OpenStudy (anonymous):

Can you explain the step?

hartnn (hartnn):

ok, first we distribute the 4 inside the brackets...